# The Multiband Imaging Survey for High-Alpha PlanetS (MISHAPS) I: Preliminary Constraints on the Occurrence Rate of Hot Jupiters in 47 Tucanae

### v1

@@ 3.2 Reference Image Photometry & Calibration
We used DoPHOT (Schechter et al., 1993; Alonso-García et al., 2012) to find stars and extract PSF fitting photometry in each $r$ and $z$ reference image. For each star, this algorithm measures pixel positions, an instrumental magnitude and uncertainty, a local sky estimate, and a fit type code indicating the type of PSF fit performed. We use this code as a data quality indicator. An astrometric solution for each DoPHOT catalog was found by matching to Gaia DR3’s 2016.0 epoch positions. Then $r$ and $z$ catalogs were merged to produce an instrumental 2-filter catalog we label “cmd.”
[…]
@@
In addition to foreground Milky Way stars, 47 Tuc’s visual proximity to the SMC and five other star clusters within it (NGC 121, ESO 28-19, ESO 28-22, HW 5, and Kron 11) requires us to carefully identify true cluster members. We crossmatch the sources identified in our images with those in Gaia DR3 (Gaia Collaboration et al., 2016; Babusiaux et al., 2023), and plot the corresponding proper motions for all sources with matches in Figure 1. We make an initial wide selection of potential cluster members, using the mean proper motion values for 47 Tuc in Baumgardt et al. (2019), $\bar{\mu}_{\rm\alpha}=5.25$ mas/yr and $\bar{\mu}_{\rm\delta}=-2.53$ mas/yr and a selection radius of
@@
where $\sigma$=0.94. Using the initial selection, we compute new values of $\bar{\mu}_{\rm\alpha}$, $\bar{\mu}_{\rm\delta}$, and their corresponding errors by averaging over the Gaia DR3 values for the selection. We make an elliptical selection on ($\bar{\mu}_{\rm\alpha},\bar{\mu}_{\rm\delta}$) and ($\sigma_{\rm\alpha},\sigma_{\rm\delta}$), recompute our values, and make another elliptical cut. To further constrain the target sample to 47 Tuc stars, we estimate the main sequence locus by plotting the color-magnitude diagram (CMD) for the proper motion selection, computing the median ($r-z$) color in 0.5 mag bins of $r$. To the binned median colors, we fit a cubic spline using the scipy.interpolate package and select stars within $\pm$0.05 mag of this fit as our final selection. This selection leaves us with 19,930 likely cluster members. The CMD, with 47 Tuc and SMC populations highlighted, is shown in Figure 2. An overview of these cuts and the resultant numbers of stars can be found in Table 2. Requiring Gaia DR3 matches effectively reduces our faint magnitude limit to $r\sim 21.0$, with significant incompleteness below $r\sim 20.0$. We note that — although we apply the proper motion cut before the injections and search are run — we do not apply the color cut until after the initial round of vetting has been performed. This choice allows us to detect EBs with CMD positions on the binary sequence which would place them outside the range of the color selection.
@@
Figure 2: $r$ vs. ($r-z$) color-magnitude diagram of our field, with only the DoPHOT type cut applied (black). Our final selection of stars is overplotted in red, and the proper motion-selected SMC stars (selected using the same procedure as for 47 Tuc stars) are overplotted in gold. The remaining black points are targets which either lie outside the 3-$\sigma$ 47 Tuc and SMC proper motion selections, or have no Gaia matches.
@@
Note. — Number of stars remaining in our sample after each cut. Our two major limitations are requiring both filters to have DoPHOT type flags of 1, and the requirement that each target have a match with Gaia for proper motion measurements. The Gaia requirement in particular cuts out many of our fainter targets, since its current magnitude limit is $G$=20.7 (corresponding to $r\approx 21.0$ for our DECam data).
@@ 3.4.1 Extinction Estimates
To determine the extinctions, $A_{\lambda}$, of each star, we first obtain estimates of $E(B-V)$ using our cross-match with the NSC, which obtains its $E(B-V)$ values from the SFD maps (Schlegel et al., 1998). However, the catalog only contains $E(B-V)$ estimates for $\sim 20\%$ of our targets. For those without an estimate, we average the $E(B-V)$ values of stars within 15” which do, and use those averages as their $E(B-V)$. We then use the $E(B-V)$ values, the central wavelengths for the DECam filters, and the Cardelli et al. (1989) extinction law, $R_{\rm V}$=3.1, to calculate $A_{r}$ and $A_{z}$ for our sample, with $A_{r}=0.083\pm 0.0046$ and $A_{z}=0.045\pm 0.0025$ being the mean extinctions.
@@ 3.4.2 Transformation to PanSTARRS System
After estimating the transformation functions, we transform our dereddened, NSC-calibrated MISHAPS data to PanSTARRS using the corresponding NSC to PS1 functions. We use the resulting $r_{\rm PS1}$ and $z_{\rm PS1}$ values in the $(r-z)_{\rm PS1}$ to $(g-i)_{\rm PS1}$ and $(i-z)_{\rm PS1}$ functions to obtain our $(g-i)_{\rm PS1}$ and $(i-z)_{\rm PS1}$ colors. Finally, we add our $z_{\rm PS1}$ to $(i-z)_{\rm PS1}$ to obtain $i_{\rm PS1}$. The spline parameters for our fit functions are given in Table 3.4.2, and the functions and their residuals are plotted along with the stars used to determine them in Figure 3. All transforms have residuals better than ±$\pm 0.03$ mag except the ($r-z$)PS1 to ($g-i$)PS1, which has residuals within ±$\pm 0.08$ mag. However, even in the case of this transform, the median residual as a function of ($r-z$) remains close to 0.
@@ 3.4.3 Radius Estimation
To estimate the radii of our stars, we use a color-diameter relation to estimate our stars’ angular diameters (e.g., Boyajian et al., 2014):
@@
where $d$ is the estimated distance to 47 Tuc (4.45 kpc, Chen et al., 2018). We use the relationship obtained in Zang et al. (2018)
@@
to determine $\log\theta_{m=0}$ This relation has typical fractional uncertainties of 3.5-6.0%.
We use the dereddened and transformed PanSTARRS magnitudes obtained in §3.4.2 to estimate our target radii, with the results shown in Figure 4. To check the derived radii, we use a MIST isochrone (Dotter, 2016; Choi et al., 2016) generated for the PanSTARRS photometric system with [Fe/H]=-0.78 (Forbes & Bridges, 2010), convert the isochrone’s absolute magnitudes to apparent magnitudes using the 4.45 kpc distance (Chen et al., 2018), and estimate the stellar radii using the luminosity and Teff. We compare the radii obtained for the isochrone stars to the radii obtained for the target stars as a function of $r$ to make sure our results are reasonable. Figure 4 shows the radii obtained for the full field as well as the isochrone, with the radii of our final stellar selection highlighted in red. The residuals between the isochrone radii and our estimated radii are shown in the bottom panel. We show only the residuals for our 19,930-star selection. The radii estimated for the rest of the stars are less likely to be accurate, since determining their radii is dependent on constraining their distances (equation 3), and thus knowing which population they belong to.
@@ 4.1 Difference Image Photometry
We use a modified version of the ISIS package (Alard & Lupton, 1998; Alard, 2000) which replaces the ISIS registration task with one based on Source-Extractor output (Bertin & Arnouts, 1996; Siverd et al., 2012) to perform difference imaging and extract difference imaging analysis (DIA) photometry. The software aligns each image, then convolves the reference image with a PSF of the spatially-varying multi-Gaussian kernel fit to match the reference image to each target image. It then subtracts the convolved reference images from the corresponding target images. Finally, it produces a photometric table of each source it extracts from the image.
@@
is the median difference between the ISIS instrumental magnitude of the reference stars $-2.5\log F_{\rm ref}$ and the DoPHOT instrumental magnitudes $m_{\rm inst}$. The median is computed using bright stars (PS1 calibrated magnitudes between $16.5$ and $18.5$ for each filter, and DoPHOT objtype flag of 1) after 5 iterations of clipping of points further than 2 standard deviations from the median to minimize the impact of outliers due to blending. We optimized the configuration parameters for ISIS by running it repeatedly on a subset of images of a field in the Galactic bulge while varying the values of one parameter at a time. We found the most crucial parameters to adjust to be rad_phot and rad_aper. These parameters represent, respectively, the radius around the target center in which the pixels will be used to estimate flux, and the radius used to normalize the flux. We found that rad_phot=5.0 px and rad_aper=6.0 px gave us the best results. Once all the relevant parameters were optimized, we ran the difference imaging process for each DECam chip.
@@ 4.2 De-Trending
Once the difference image analysis (DIA) photometry has been extracted and calibrated, systematic variations are removed using the Trend-Filtering Algorithm (TFA) implemented through VARTOOLS (Kovács et al., 2005). This process is performed chip-by-chip, as each of DECam’s CCDs has slightly different systematics. Usage of TFA involves selecting a subset of lightcurves with low variation — referred to as “trend stars” —which are used to determine the systematic trends in the data. To choose the size of this subset, we ran the TFA on lightcurves from the N10 chip multiple times, varying the number of trend stars used in each run. We then defined a “quasi-reduced $\chi^{2}$” metric
[…]
@@
where $\log(RMS)$ is the logarithm of each measurement’s RMS, $\rho$ is the median of $\log(RMS)$, and MAD($\sigma_{R}$) is the median absolute deviation of the scatter scaled to equal $\sigma$ for Gaussian noise. The remaining stars were then randomly sampled for 15 stars. To check the resultant photometric precision, we plot the MAD for each target as a function of $r-$magnitude and compare it to the estimated transit depths for planets of various radii (Figure 5).
@@ 5.1 Transit Injection
To fully characterize the efficiency of our search and calculate an occurrence rate, we need to inject a range of synthetic transits into our lightcurves and analyze how well our search recovers them. Using a sample of 19,930 cluster stars with radius estimates, we use the JKTEBOP code (Southworth, 2013) to calculate transit models and inject them into our data.
@@
Orbital periods, $P$, between 0.5–10.0 days and planet radii, $R_{\rm p}$, between 0.5–2.0 $R_{\rm jup}$ were drawn from uniform distributions for each system. These limits are selected to encompass the ranges considered by G00 and MW17, as well as the ranges considered for Kepler stars in Fressin et al. (2013), allowing us to more carefully compare our occurrence rate results with those of the Milky Way disk. Inclinations were randomly selected from the range of $\cos i$ angles which would allow a full or grazing transit. The maximum possible $\cos i$ value is also the transit probability for an isotropic distribution of circular orbits,
@@
We injected the resulting JKTEBOP models into our de-trended lightcurves. [^2] Future iterations will inject the transits into the original lightcurves, and the injections will be de-trended along with the real lightcurves. A limited sample ($\sim$2,000 injections) was run and the lightcurves were plotted with parameters and ID number visible to determine whether we were generating models with realistic radii and with depths, limb-darkening effects, and transit times that matched what is expected for systems with similar characteristics. These injections were not used in our detection efficiency determination, but the original lightcurves were recycled to create new injections once we were confident in our models. We created $\sim$40,000 injection lightcurves from $\sim$20,000 unique lightcurves, by injecting a transiting planet into each original lightcurve twice. We chose the total number of injections to achieve binomial uncertainties on the detection efficiency $\sim$5in bins of $P$ and $R_{\rm p}$. A single lightcurve example with multiple full and partial transits is shown in Figure 6. Grids of transit injections with varying $P$, $R_{\rm p}$, and $r-$magnitude can be found in Figures 19-21 of Appendix A.
Figure 6: Example of injected full and partial transits of a $1.26R_{\rm Jup}$ planet orbiting a $0.59R_{\rm\odot}$, $r=19.9$ star in our data with $P$=0.97 days. Each panel shows a different night, with the top two panels showing nights flagged as having full transits, and the bottom two showing nights flagged as partial. The $r$ band data are shown in blue, and the $z$ is shown in red. The black line is the injected model, and the green line is the best boxcar model. We note that the top left transit is actually a partial transit since we do not capture the full egress in the data, but the algorithm flags it as a full transit based on its best S/N, $t_{c}$, and $\tau$. Our detection algorithm was able to successfully recover both the full and partial transits with S/N $>$7.

### v2

@@ 3.2 Reference Image Photometry & Calibration
We used DoPHOT (P. L. Schechter et al., 1993; J. Alonso-Garc´ıa et al., 2012) to find stars and extract PSF fitting photometry in each $r$ and $z$ reference image. For each star, this algorithm measures pixel positions, an instrumental magnitude and uncertainty, a local sky estimate, and a fit type code indicating the type of PSF fit performed. We use this code as a data quality indicator. An astrometric solution for each DoPHOT catalog was found by matching to Gaia DR3’s 2016.0 epoch positions. Then $r$ and $z$ catalogs were merged to produce an instrumental 2-filter catalog we label “cmd.”
Following the method of W. Zang et al. (2018), we calibrate the DoPhot photometry to the NOIRLab Source Catalog Data Release 2 (NSC DR2 D. L. Nidever et al., 2018, 2021), fitting for a zeropoint offset in each chip independently and a color term that is common to all chips in the field. This is done for stars with $7.0\lesssim r_{\rm NSC}\lesssim 18.511$1The actual cuts on stars used for calibration are done on the instrumental magnitudes shifted to the NSC magnitude scale only by the mean difference from instrumental to NSC magnitude for all matched stars, combined reported NSC and instrumental magnitude errors $\sqrt{\sigma_{r,{\rm NSC}}^{2}+\sigma_{r,{\rm inst}}^{2}}<0.01$, and instrumental color $(r-z)_{\rm inst}$ color between 0.1 mag less than the mean instrumental color and 0.8 mag greater than the mean instrumental color. This selection includes the 47 Tuc turnoff, much of the SMC giant branch above the clump, and a scattering of field stars. In sparse chips, the median offset between our calibrated magnitude and the NSC DR2 magnitude is typically $\lesssim 5$ mmag, and the standard deviation is typically between 0.01 and 0.02 mag (estimated using a median absolute deviation scaled by 1.48). The densest chips we analyze in this work, N10 and S10, have median offsets of 0.013 and 0.014 mag, respectively, and standard deviations of 0.039 and 0.038 mag, respectively. The worse performance of the photometric calibration in more crowded fields is a result of the aperture photometry used in NSC DR2 photometry, which is more affected by blending than our PSF photometry.
@@
In addition to foreground Milky Way stars, 47 Tuc’s visual proximity to the SMC and five other star clusters within it (NGC 121, ESO 28-19, ESO 28-22, HW 5, and Kron 11) requires us to carefully identify true cluster members. We crossmatch the sources identified in our images with those in Gaia DR3 (Gaia Collaboration et al., 2016; C. Babusiaux et al., 2023), and plot the corresponding proper motions for all sources with matches in Figure 1. We make an initial wide selection of potential cluster members, using the mean proper motion values for 47 Tuc in H. Baumgardt et al. (2019), $\bar{\mu}_{\rm\alpha}=5.25$ mas/yr and $\bar{\mu}_{\rm\delta}=-2.53$ mas/yr and a selection radius of
@@
where $\sigma$=0.94. Using the initial selection, we compute new values of $\bar{\mu}_{\rm\alpha}$, $\bar{\mu}_{\rm\delta}$, and their corresponding errors by averaging over the Gaia DR3 values for the selection. We make an elliptical selection on ($\bar{\mu}_{\rm\alpha},\bar{\mu}_{\rm\delta}$) and ($\sigma_{\rm\alpha},\sigma_{\rm\delta}$), recompute our values, and make another elliptical cut. To further constrain the target sample to 47 Tuc stars, we estimate the main sequence locus by plotting the color-magnitude diagram (CMD) for the proper motion selection, computing the median (rzr-z) color in 0.5 mag bins of $r$. To the binned median colors, we fit a cubic spline using the scipy.interpolate package and select stars within $\pm$0.05 mag of this fit as our final selection. This selection leaves us with 19,930 likely cluster members. The CMD, with 47 Tuc and SMC populations highlighted, is shown in Figure 2. An overview of these cuts and the resultant numbers of stars can be found in Table 2. Requiring Gaia DR3 matches effectively reduces our faint magnitude limit to $r\sim 21.0$, with significant incompleteness below $r\sim 20.0$. We note that — although we apply the proper motion cut before the injections and search are run — we do not apply the color cut until after the initial round of vetting has been performed. This choice allows us to detect EBs with CMD positions on the binary sequence which would place them outside the range of the color selection.
@@
Figure 2: $r$ vs. (rzr-z) color-magnitude diagram of our field, with only the DoPHOT type cut applied (black). Our final selection of stars is overplotted in red, and the proper motion-selected SMC stars (selected using the same procedure as for 47 Tuc stars) are overplotted in gold. The remaining black points are targets which either lie outside the 3-$\sigma$ 47 Tuc and SMC proper motion selections, or have no Gaia matches.
@@
Note. — Number of stars remaining in our sample after each cut. Our two major limitations are requiring both filters to have DoPHOT type flags of 1, and the requirement that each target have a match with Gaia for proper motion measurements. The Gaia requirement in particular cuts out many of our fainter targets, since its current magnitude limit is GG=20.7 (corresponding to $r\approx 21.0$ for our DECam data).
@@ 3.4.1 Extinction Estimates
To determine the extinctions, $A_{\lambda}$, of each star, we first obtain estimates of E(BV)E(B-V) using our cross-match with the NSC, which obtains its E(BV)E(B-V) values from the SFD maps (D. J. Schlegel et al., 1998). However, the catalog only contains E(BV)E(B-V) estimates for $\sim 20\%$ of our targets. For those without an estimate, we average the E(BV)E(B-V) values of stars within 15” which do, and use those averages as their E(BV)E(B-V). We then use the E(BV)E(B-V) values, the central wavelengths for the DECam filters, and the J. A. Cardelli et al. (1989) extinction law, $R_{\rm V}$=3.1, to calculate $A_{r}$ and $A_{z}$ for our sample, with $A_{r}=0.083\pm 0.0046$ and $A_{z}=0.045\pm 0.0025$ being the mean extinctions.
@@ 3.4.2 Transformation to PanSTARRS System
After estimating the transformation functions, we transform our dereddened, NSC-calibrated MISHAPS data to PanSTARRS using the corresponding NSC to PS1 functions. We use the resulting $r_{\rm PS1}$ and $z_{\rm PS1}$ values in the $(r-z)_{\rm PS1}$ to $(g-i)_{\rm PS1}$ and $(i-z)_{\rm PS1}$ functions to obtain our $(g-i)_{\rm PS1}$ and $(i-z)_{\rm PS1}$ colors. Finally, we add our $z_{\rm PS1}$ to $(i-z)_{\rm PS1}$ to obtain $i_{\rm PS1}$. The spline parameters for our fit functions are given in Table 3.4.2, and the functions and their residuals are plotted along with the stars used to determine them in Figure 3. All transforms have residuals better than $\pm 0.03$ mag except the (rzr-z)PS1 to (gig-i)PS1, which has residuals within $\pm 0.08$ mag. However, even in the case of this transform, the median residual as a function of (rzr-z) remains close to 0.
@@ 3.4.3 Radius Estimation
To estimate the radii of our stars, we use a color-diameter relation to estimate our stars’ angular diameters (e.g., T. S. Boyajian et al., 2014):
@@
where $d$ is the estimated distance to 47 Tuc (4.45 kpc, S. Chen et al., 2018). We use the relationship obtained in W. Zang et al. (2018)
@@
to determine $\log\theta_{m=0}$ . This relation has typical fractional uncertainties of 3.5-6.0%.
We use the dereddened and transformed PanSTARRS magnitudes obtained in §3.4.2 to estimate our target radii, with the results shown in Figure 4. To check the derived radii, we use a MIST isochrone (A. Dotter, 2016; J. Choi et al., 2016) generated for the PanSTARRS photometric system with [Fe/H]=-0.78 (D. A. Forbes & T. Bridges, 2010), convert the isochrone’s absolute magnitudes to apparent magnitudes using the 4.45 kpc distance (S. Chen et al., 2018), and estimate the stellar radii using the luminosity and Teff. We compare the radii obtained for the isochrone stars to the radii obtained for the target stars as a function of $r$ to make sure our results are reasonable. Figure 4 shows the radii obtained for the full field as well as the isochrone, with the radii of our final stellar selection highlighted in red. The residuals between the isochrone radii and our estimated radii are shown in the bottom panel. We show only the residuals for our 19,930-star selection. The radii estimated for the rest of the stars are less likely to be accurate, since determining their radii is dependent on constraining their distances (equation 3), and thus knowing which population they belong to.
@@ 4.1 Difference Image Photometry
We use a modified version of the ISIS package (C. Alard & R. H. Lupton, 1998; C. Alard, 2000) which replaces the ISIS registration task with one based on Source-Extractor output (E. Bertin & S. Arnouts, 1996; R. J. Siverd et al., 2012) to perform difference imaging and extract difference imaging analysis (DIA) photometry. The software aligns each image, then convolves the reference image with a PSF of the spatially-varying multi-Gaussian kernel fit to match the reference image to each target image. It then subtracts the convolved reference images from the corresponding target images. Finally, it produces a photometric table of each source it extracts from the image.
@@
is the median difference between the ISIS instrumental magnitude of the reference stars $2.5\log F_{\rm ref}$ and the DoPHOT instrumental magnitudes $m_{\rm inst}$. The median is computed using bright stars (PS1 calibrated magnitudes between $16.5$ and $18.5$ for each filter, and DoPHOT objtype flag of 1) after 5 iterations of clipping of points further than 2 standard deviations from the median to minimize the impact of outliers due to blending. We optimized the configuration parameters for ISIS by running it repeatedly on a subset of images of a field in the Galactic bulge while varying the values of one parameter at a time. We found the most crucial parameters to adjust to be rad_phot and rad_aper. These parameters represent, respectively, the radius around the target center in which the pixels will be used to estimate flux, and the radius used to normalize the flux. We found that rad_phot=5.0 px and rad_aper=6.0 px gave us the best results. Once all the relevant parameters were optimized, we ran the difference imaging process for each DECam chip.
@@ 4.2 De-Trending
Once the difference image analysis (DIA) photometry has been extracted and calibrated, systematic variations are removed using the Trend-Filtering Algorithm (TFA) implemented through VARTOOLS (G. Kovács et al., 2005). This process is performed chip-by-chip, as each of DECam’s CCDs has slightly different systematics. Usage of TFA involves selecting a subset of lightcurves with low variation — referred to as “trend stars” —which are used to determine the systematic trends in the data. To choose the size of this subset, we ran the TFA on lightcurves from the N10 chip multiple times, varying the number of trend stars used in each run. We then defined a “quasi-reduced $\chi^{2}$” metric
$\text{quasi-}\chi_{\rm red}^{2}=\frac{1}{1.085^{2}}\frac{\sum_{i}^{N_{\rm data}}(RMS_{i}^{2}/\sigma_{i}^{2})}{N_{\rm data}-(N_{\rm trend}+1)}$ (7)
@@
where $\log(RMS)$ is the logarithm of each measurement’s RMS, $\rho$ is the median of $\log(RMS)$, and MAD($\sigma_{R}$) is the median absolute deviation of the scatter scaled to equal $\sigma$ for Gaussian noise. From the stars that survive these cuts, we randomly select 15 to serve as our trend stars for TFA. To check the resultant photometric precision, we plot the MAD for each target as a function of $r$-magnitude and compare it to the estimated transit depths for planets of various radii (Figure 5).
@@ 5.1 Transit Injection
To fully characterize the efficiency of our search and calculate an occurrence rate, we need to inject a range of synthetic transits into our lightcurves and analyze how well our search recovers them. Using a sample of 19,930 cluster stars with radius estimates, we use the JKTEBOP code (J. Southworth, 2013) to calculate transit models and inject them into our data.
@@
Orbital periods, $P$, between 0.5–10.0 days and planet radii, $R_{\rm p}$, between 0.5–2.0 $R_{\rm jup}$ were drawn from uniform distributions for each system. These limits are selected to encompass the ranges considered by G00 and MW17, as well as the ranges considered for Kepler stars in F. Fressin et al. (2013), allowing us to more carefully compare our occurrence rate results with those of the Milky Way disk. Inclinations were randomly selected from the range of $\cos i$ angles which would allow a full or grazing transit. The maximum possible $\cos i$ value is also the transit probability for an isotropic distribution of circular orbits,
@@
We injected the resulting JKTEBOP models into our de-trended lightcurves. [^2] Future iterations will inject the transits into the original lightcurves, and the injections will be de-trended along with the real lightcurves. A limited sample ($\sim$2,000 injections) was run and the lightcurves were plotted with parameters and ID number visible to determine whether we were generating models with realistic radii and with depths, limb-darkening effects, and transit times that matched what is expected for systems with similar characteristics. These injections were not used in our detection efficiency determination, but the original lightcurves were recycled to create new injections once we were confident in our models. We created $\sim$40,000 injection lightcurves from $\sim$20,000 unique lightcurves, by injecting a transiting planet into each original lightcurve twice. We chose the total number of injections to achieve binomial uncertainties on the detection efficiency $\sim$5% in bins of $P$ and $R_{\rm p}$. A single lightcurve example with multiple full and partial transits is shown in Figure 6. Grids of transit injections with varying $P$, $R_{\rm p}$, and $r$-magnitude can be found in Figures 19-21 of Appendix A.
Figure 6: Example of injected full and partial transits of a $1.26R_{\rm Jup}$ planet orbiting a $0.59R_{\rm\odot}$, $r=19.9$ star in our data with PP=0.97 days. Each panel shows a different night, with the top two panels showing nights flagged as having full transits, and the bottom two showing nights flagged as partial. The $r$ band data are shown in blue, and the $z$ is shown in red. The black line is the injected model, and the green line is the best boxcar model. We note that the top left transit is actually a partial transit since we do not capture the full egress in the data, but the algorithm flags it as a full transit based on its best S/N, $t_{c}$, and $\tau$. Our detection algorithm was able to successfully recover both the full and partial transits with S/N >>7.